OpenStudy (anonymous):

Hi, still having problems understanding how to add 5/8 + 1/4 pls help

OpenStudy (anonymous):

You first want to find a common denominator. In this case that common denomiator is 8

OpenStudy (anonymous):

so you should have : 5/8 +2/8

OpenStudy (anonymous):

Now you just add the numerators to get 7/8

OpenStudy (anonymous):

Make sense?

OpenStudy (idily101):

that is correct actually

OpenStudy (anonymous):

DO you understand how to find a common denominator?

OpenStudy (idily101):

me or somebody else

OpenStudy (anonymous):

sorry I was talking to shiloe

OpenStudy (anonymous):

Where is the 2/8 coming from when the problem says 1/4 no I understand that 2x1=2 and 2x4=8 hence 2/8 but I don't understand how to find the common denominator

OpenStudy (anonymous):

Your common denominator is 8 so basically what you are doing is having the fractions have the same denominator but in order to do that you have to multiply both the top and bottom of 1/4 by 2 to get 2/8 just like you did above. 2/8 =1/4 same fraction just larger numbers... make more sense?

OpenStudy (anonymous):

Thank you for the help. I'm still not sure, but I'll use your guideline and perhaps I'll get it with practice. It's been soooo long since I've done these that I've forgotten. Thanks again

OpenStudy (anonymous):

Ha ha no problem If you have another problem like that one I would be more than willing to go through it again and try to make it easier to understand

OpenStudy (anonymous):

you get 8 because currently 8 is the highest denominator and 4 goes evenly into it so you would choose to use that number so that you can add the two parts of the problem together a little easier Say for example you have 2 pies. You cut one of them into 4 pieces and one into 3 pieces.. it would be difficult to give the slices to your friends because the one that was cut into 3 pieces would give bigger pieces so in the problem we are just changing the 4 into an 8 so that we can add the two parts together.